Session: Feb 2005 (session 1)
Wavelength: 1.3 cm


To give feedback on this experiment, click here

Experiment Feedback

Urumqi (yang wen jun at Sun Feb 20 11:05:51 2005) Minor failures :
It is good weather.The cable-cal detector doesn't work in Urumqi,and P-cal detector is broken-down, the communication between the fs and antenna computer had some trouble during the experiment,so lost few data,more details explain in log file.

Medicina (Sergio Poppi at Sun Feb 20 11:15:43 2005) Success :
No problems during the experiment. Rainy weather at Medicina, from 20 feb. morning.

Onsala (Michael Lindvqvist at Sun Feb 20 14:50:45 2005) Success :
Antenna in standby UT 02:25-02:26. Otherwise no known problems.

Shanghai (Tao An at Mon Feb 21 01:30:46 2005) Success :

Effelsberg (A.Kraus at Mon Feb 21 12:34:56 2005) Minor failures : WIND
Various stops due to snowfall: UT1112-1152, UT1445-1540, UT1620-1901, and UT1920-0134.

Metsahovi (Ari Mujunen at Mon Feb 21 13:15:45 2005) Minor failures : LATE
The schedule was set up for old VLBA rack, although we currently have a VLBA4. This was noticed after three first passes were (not) recorded in VLBA mode. Re-drudged changing equipment to VLBA4 and restarted, apparently ok for the rest of the schedule. Pcal turned manually on, cable cal measurement system on, prints out non-standard log lines but the numbers are there.

Jodrell_Bank (Alastair Gunn at Mon Feb 21 21:06:32 2005) Minor failures :
MkII was used. Few minutes off source at about 10:29 UT due to elevation drive fault. No other known problems.

Cambridge (Alastair Gunn at Mon Feb 21 21:07:08 2005) Success :
No known problems.

Noto (P.Cassaro at Sat Feb 26 16:52:06 2005) Success :
No known problems

MPIfR_Bonn (Walter Alef at Wed Mar 30 12:41:47 2005) Minor failures :
* Ef: recorded with disc. Top 2 channels have poor statistics (as expected), needs correction with sampler statistics. Pcal weak, higher-freq pcals tried but no use. Need manual PCals, rather hard to determine. * Eb: recorded parallel on tape. Used different track pattern to standard. had to adjust bit density to 56700 to get 160ips. Pcal weak, need manual.Statistics better than Ef but tape decode not perfect (try changing tape unit). * Jb: top 4 channels useless. These should have been on high patch but were obviously not. Duhh... * Cm: only ONE frequency channel (the first). * Mh: No fringe. 1 Disc in pack is bad. No phase-cal, bandpass suspiciously flat. * Nt: Plays OK but no fringe. Bandpass is entirely flat, no sign of filter shape. No phase cal, so obviously not connected to sky. * Sh: No fringe, tried 2 scans and wide delay window, +- 1 sec. Saw SNR 7.1 with -1sec, but probably not real. Diagnosis unclear. * Mc,On,Ur: OK (Total 6 useful stations)